Melinda invests $410,000 in a City of Heflin bond that pays 6.4 percent interest. Alternatively, Melinda could have invested the $410,000 in a bond recently issued by Surething Inc., that pays 8 percent interest and has risk and other nontax characteristics similar to the City of Heflin bond. Assume Melinda’s marginal tax rate is 20 percent. (Leave no cells blank - be sure to enter "0" wherever required. Round your after-tax rate of return to one decimal place.) Required: How much explicit tax does Melinda pay on the City of Heflin bond? How much implicit tax does she pay on the City of Heflin bond? How much explicit tax would she have paid on the Surething Inc. bond? What is her after-tax rate of return on the Surething Inc. bond?
Answer #1
Details Provided
Melindas Investment 410,000
Interest 6.40%
Melinda could have invested in bond by SureThing Inc 410,000
Interest 8%
Melinda’s marginal tax rate 20%
1). How much explicit tax does Melinda pay on the City of Heflin bond?
Heflin bond is a tax exempt bond. So she pays no explicit tax on the interest from City of Hefin bond
2). How much implicit tax does she pay on the City of Heflin bond?
Interest on City of Heflin bond 410,000*6.40% =            26,240 (b)
Interest on City of Surething Bond 410,000*8%=            32,800 (a)
Implicit Tax She earns 26,240 interest on City of Heflin bond. Surething's could pay her 32,800 of taxable interest. So Melinda pays 6,560 implicit tax on Cit of Heflin bond              6,560 =(a)-(b)
3). How much explicit tax would she have paid on the Surething Inc. bond?
Marginal Tax 20% (c )
Interest on City of Surething Bond            32,800 (d)
explicit tax would she have paid on the Surething Inc. =              6,560 =(c )*(d)
4). What is her after-tax rate of return on the Surething Inc. bond?
Interest Income            32,800
Tax              6,560
32800-6560=            26,240
Melindas Investment          410,000
after-tax rate of return on the Surething Inc. bond 26240/410000= 6.4%